Can you help me find the answer to this math question? Three consecutive odd integers such that three times the smallest is nine more than twice the largest. Find the largest integer.
Answer by mananth(16949)

let the three consequtive odd integers be x, x+2,x+4
3 times the smallest = 3x
2 times the largest = 2(x+4)
9 more
2(x+4) +9
3x=2(x+4)+9
3x=2x+8 +9
3x-2x= 17
x= 13
17 ,19,21 are the integers